### v2

# Generalizable Graph Neural Networks for Robust Power Grid Topology Control

Abstract
The energy transition necessitates new congestion management methods. One such method is controlling the grid topology with machine learning (ML). This approach has gained popularity following the Learning to Run a Power Network (L2RPN) competitions. Graph neural networks (GNNs) are a class of ML models that reflect graph structure in their computation, which makes them suitable for power grid modeling. Various GNN approaches for topology control have thus been proposed. We propose the first GNN model for grid topology control that uses only GNN layers. Additionally, we identify the busbar information asymmetry problem that the popular homogeneous graph representation suffers from, and propose a heterogeneous graph representation to resolve it. We train both homogeneous and heterogeneous GNNs and fully connected neural networks (FCNN) baselines on an imitation learning task. We evaluate the models according to their classification accuracy and grid operation ability. We find that the heterogeneous GNNs perform best on in-distribution networks, followed by the FCNNs, and lastly, the homogeneous GNNs. We also find that both GNN types generalize better to out-of-distribution networks than FCNNs.
keywords:
Graph Neural Networks , Topology Control , Heterogeneous Graph Representation , Busbar Information Asymmetry , L2RPN , Out-of-Distribution Generalization
journal: Applied Energy
[…]
{graphicalabstract}
{highlights}
Heterogeneous over homogeneous graphs increase graph neural network expressiveness
Heterogeneous graph neural networks are the highest-performing models
Graph neural networks generalize better to out-of-distribution networks
Machine learning agents obtain good performance at high efficiency
1 Introduction
The energy transition is crucial for ensuring society’s sustainability and future security. Transmission system operators (TSOs) play a crucial role in this transition. They are consequently facing new operational challenges [9]. Grid congestion is a problem that TSOs are already encountering but which the energy transition will exacerbate [42]. If not mitigated, grid congestion could make the future availability of energy unreliable.
There are several actions that TSOs can take to alleviate grid congestion [17]. Presently, redispatching and curtailment are the primary methods. However, these actions are often costly, disruptive, or not applicable. Topology control presents an opportunity for remedying grid congestion without these problems [41]. Congestion can be relieved by exploiting the flexibility in the network topology. However, this technique mostly remained unutilized, as the space of potential topologies is combinatorially large.
Due to the recent congestion management challenges, there has been a new interest in topology control. This interest led to the development of the GridOptions tool for remedial action recommendation by Viebahn et al. [41]. This tool validates the practical value of topology control. Even simple topological strategies could reduce line congestion by ten to twenty percent. The recommended strategies improved upon strategies without topological actions or those considered by grid operators. However, while the results are promising, further developments are necessary for broader applicability. The tool was only applied to a small subset of the Dutch power grid, featuring nine substations. Additionally, the inference time of the tool is relatively long: one hour. The next challenges involve scaling this approach and reducing inference time.
Machine learning (ML) presents a solution to these challenges [42]. ML models can learn which topologies are suitable in which situations and predict these topologies much quicker than traditional computational methods. To facilitate research in this area, the Learning to Run a Power Network (L2RPN) competitions were hosted [28, 27]. Furthermore, the Grid2Op Python library was developed. [^1] https://grid2op.readthedocs.io/ This library greatly simplified the development and evaluation of ML methods for grid topology control. These developments kick-started research into ML methods for grid topology control.
Classical ML models are typically designed for tabular data and hence fail to capture the graph structure of power grids. This is problematic as the grid topology frequently changes, e.g., because of maintainance. Power grid operation requires adapting to such changes. Graph neural networks (GNNs) [36] are a class of ML models for graph-structured data. The computation in GNNs can reflect arbitrary graphs, which allows them to exploit a graph’s structure. This makes GNNs suitable for modeling power grids. GNNs have thus been applied to many grid operation tasks [5, 4, 31, 11, 18, 13, 30].
Various GNN approaches have been proposed for grid topology control [45, 43, 44, 35, 40, 33]. This research has demonstrated the feasibility of GNN approaches for grid topology control. However, important topics have not yet been addressed:
  1. 1.
    All approaches that directly predict topology actions with GNNs only use GNN layers as feature extractors or embedders [45, 43, 44, 33]. Fully-connected or attention layers follow the GNN layers. No research has investigated ’fully graphical’ neural networks for grid topology control. Purely GNN approaches might be able to exploit the topology more deeply.
  2. 2.
    All present literature uses the graph representation specified by which busbars and power lines objects share. However, this is only one of many ways to represent the power grid as a graph. There has not been any study that compares graph representations and their effect on GNN performance.
  3. 3.
    Grid changes are frequent, and operation needs to continue despite these changes. For robust operation, models must generalize to potentially unexpected network states. A previous competition has focused on generalization to out-of-distribution injection profiles [27]. However, no work has investigated generalization to out-of-distribution networks.
This study aims to fill the aforementioned gaps in applying GNNs to grid topology control. We develop fully GNN models that use a richer graph representation. The main contributions of this study are:
  1. 1.
    We propose the first fully graphical neural network approach for grid topology control. The models are trained on a multi-label binary classification imitation learning task involving various networks and topologies. We compare the GNNs against fully connected neural networks (FCNNs). We evaluate the models by their classification accuracy and ability to operate the power grid as agents. We also consider expert and hybrid agents and compare their inference speed.
  2. 2.
    We first analyze the effect of grid graph representation on GNN performance. The graph representation in the literature models current but not potential object-busbar attachments. This results in busbar information asymmetry, which blocks information flow and impairs GNN expressiveness. We propose a heterogeneous graph representation that models current and potential connections, which remedies busbar information asymmetry. We evaluate both graph representations with GNNs.
  3. 3.
    We first investigate the ability of graph neural networks to generalize topology control to out-of-distribution grids. Generalization is evaluated w.r.t. model accuracy and the ability of agents to operate the power grid, and contrasted against FCNNs.
This paper is organized as follows. Section 2 describes related work in ML grid topology control. Section 3 describes the power grid setup used in this study. Section 4 describes the methods used to develop the ML models. Section 5 describes the evaluation and analysis of the ML models. Sections 6 and 7 provide discussion and recommendations for future work, respectively. The code for this project is available. [^2] https://github.com/MatthijsdeJ/GNN_PN_Imitation_Learning Table 1 lists the meaning of mathematical symbols used in this paper. Table 2 lists the glossary.
Table 1: Mathematical symbols used in this paper and their meaning.
Notation Meaning
$O$ The set of grid objects (i.e., generators, loads, and line endpoints).
$S$ The set of substations.
$\textbf{x}_{u}$ The features of a GNN model at node $u\in O$.
$\textbf{h}_{u,k}$ The node embeddings of a GNN model at node $u\in O$ in layer $k$.
$\mathcal{N}(u)$ The neighbors of node $u\in O$.
$W$ A weight matrix.
$\sigma$ An activation function.
$\textbf{p}=(0,1)^{|O|}$ The output of a ML model.
$\textbf{y}=\{0,1\}^{|O|}$ The target of a ML model.
$\alpha=0.1$ The label weight hyperparameter.
$\textbf{w}=\{\alpha,1\}^{|O|}$ The label weights for the loss.
$\{\textbf{p}_{s}\}{}\forall s\in S$ A partition of the output vector p into subvectors $\textbf{p}_{s}$. Subvector $\textbf{p}_{s}$ corresponds to substation $s\in S$.
$\eta$ = 0.97 The activity threshold parameter.
$\theta=1.0$ The risk threshold parameter, used by various agents.
Table 2: Glossary
Term Meaning
Busbar Information Asymmetry An asymmetry in the availability of information of objects on busbars.
Full-network Regime An environmental regime without statically or randomly disabled lines.
In-distribution Data (ID) Data with networks that the models are exposed to during training.
Injection The generators and loads of the environment.
Loading The ratio of a line’s current over its thermal limit.
Multi-label Binary Classification A learning task where the target consists of multiple binary values.
N-1 Network A variation of the default network with a single line disabled.
N-1 Redundancy The state of the network being resilient to the disablement of any single line.
Object Generators, loads, and line endpoints in the Grid2Op environment.
Out-of-distribution Data (OOD) Data with networks that the models are not exposed to during training; used to investigate generalization.
Planned-outage Regime An environmental regime with statically disabled lines.
Topology Reversal The periodical resetting of the topology to the default topology.
Topology Vector The Grid2Op variable that specifies the object-busbar configurations.
Unplanned-outage Regime An environmental regime with an opponent that randomly disables lines.
2 Related Work
The L2RPN competitions piqued the academic interest in grid topology control with machine learning. Prior research into topology control used linear programming [10, 49], but such methods are too rigid and/or slow to be practical. The first L2RPN competition premiered in 2019. It focused on operating scenarios of the EEE14-bus network within a specific time frame [28]. The following L2RPN competition was featured at WCCI in 2020 and featured the larger IEEE118-bus network [29]. The next competition at NeurIPS 2020 [27] introduced a robustness track with unplanned outages and an adaptability track with out-of-distribution injection profiles. Later L2RPN competitions focused on trust and sending an alarm signal [26], and the inclusion of batteries [37].
Graph neural networks have been successfully applied to many tasks of power grid operation, among which power flow estimation [5, 4], optimal power flow computation [31, 11], load shedding [18], outage prediction [13], and stability prediction [30]. For a comprehensive survey on graph reinforcement learning for power grids, we recommend the survey by Hassouna et al. [14].
Several studies investigated GNNs for topology control. Yoon et al. [45] propose a hierarchical actor-critic approach that uses a GNN module to obtain graph embeddings. Their approach won the L2RPN WCCI 2020 challenge. Xu et al. [43] are the first to apply GNNs with graph attention layers for grid topology control. They apply this GNN with a proximal policy optimization approach with an action space searched with monte-carlo tree search. Their approach can, on average, maintain the system for four days in the 2020 NeurIPS L2RPN robustness track. Xu et al. [44] find that a deep reinforcement learning agent which uses features extracted with a GNN trains more quickly and achieve better performance than an agent without. Van der Sar et al. [35] propose a hierarchical actor-critic approach with one agent per substations. These agents use graph neural networks blocks, although the details are left unspecified. They apply this approach to the IEEE5-bus network. Qiu et al. [33] propose a method similar to Yoon et al.’s [45] and obtain similar results.
Taha et al. [40] propose a different approach than directly predicting topology actions with GNNs. They use GNNs for grid estimation. These estimates are then used in grid topology tree search. We will also discuss the approach by Zhao et al. [47], which is relevant because of their focus on generalization. They propose a proximal policy optimization for redispatching that relies on graph embeddings. The graph embeddings are generated by a GraphSAGE model trained with a unsupervised learning task. They find that this graph module can succesfully represent out-of-distribution grid changes. Other 2020 NeurIPS L2RPN submissions also used graph neural networks but were not notably successful [27].
Other work did not use graph neural networks. These approaches have mostly used reinforcement learning (RL). Chauhan et al. [2] propose the PowRL framework, which integrates a novel heuristic with an RL approach. They achieved the best results on the legacy 2020 L2RPN robustness track yet. Dorfer et al. [6] adapt the AlphaZero framework [38] for the power grid domain. They achieved top results in the legacy 2020 WCCI L2RPN competition. Pan et al. [32] investigate the vulnerability of RL methods to adversarial attacks. They propose a more robust agent trained adversarially. Subramanian et al. [39] perform an in-depth investigation of a simple agent on a single scenario. Manczak et al. [24] propose a hierarchical RL approach that separates substation and configuration selection.
Imitation learning (IL) has also been applied to the problem of grid topology control. De Jong et al. [15] recently performed the first investigation in IL as a stand-alone method for topology control. They found that both IL and hybrid agents could obtain good performance and inference speed combinations. Prior work on IL used IL as a ’warm start’ for RL. Lan et al. [20] first proposed IL to pre-train an RL model for grid control. Binbinchen [8] achieved second place in the 2020 NeurIPS L2RPN submission by expanding this approach. Lehna et al. [22, 21] refined this approach. They introduced topology reversal and a higher-performing N-1 expert agent [22], and a heuristic target topology approach [21]. The winning submission of the 2023 L2RPN competition also included this approach [19].
Other approaches used neither RL nor IL. Zhou et al. [48] use evolutionary strategies to train a model to narrow down the action space. The submission that won both 2020 NeurIPS L2RPN tracks used a similar approach [27]. Although not ML, the expert system developed by Marot et al. has also achieved notoriety [25].
3 Power Grid Setup
3.1 Power Grid
In this study, we experiment on the IEEE14-bus system, displayed in Figure 1. We simulate this system with the rte_case14_realistic environment from the Grid2Op library [^3] This environment has since been deprecated. Future research should use environment l2rpn_case14_sandbox.. The power grid includes fourteen substations, five generators, 11 loads, and 20 power lines. One generator is solar, one nuclear, one wind-based, and two are thermal. The grid is divided into two sides. The high-voltage transmission side contains substations 0 to 4. The low-voltage distribution side contains substations 5 to 13. Lines 15 to 19 model the transformers connecting the two sides of the grid. We adjust the thermal limits to the values specified by Subramanian et al. [39]. This makes the transmission and distribution differences more pronounced and realistic.
We also investigate network variations with single lines disabled. We refer to these as N-1 networks. We use the N-1 networks to investigate the ability of agents to operate with line outages and to test the generalization of ML models to out-of-distribution networks.
Figure 1: The default state of Grid2Op environment rte_case14_realistic.
3.2 Operational Period
The environment features one thousand scenarios, each consisting of 8064 five-minute timesteps. One scenario thus models 28 days. Each scenario features different injection profiles. We split the scenarios into individual days. This daily period better reflects operational periods and emulates beneficial topology reversal [22]. A game-over occurs when the power grid fails to transport sufficient power from generators to loads [^4] Three scenarios were excluded where Grid2Op failed to converge the power flow without agent misoperation. These scenarios were omitted throughout the study..
3.3 Action Space
The substations have precisely two busbars in the environment. The topology actions switch objects, i.e., generators, loads, and line endpoints, between these busbars. The topology vector specifies the busbar attachments of objects. Each index in the vector corresponds to an object. At an index, a value of 1 indicates attachment to the first busbar, 2 indicates attachment to the second busbar, and -1 indicates object disconnection.
Actions are limited to a single substation per timestep to reflect real-world timing constraints. Certain substation configurations are invalid (isolated generators or loads) or redundant (configurations mirrored w.r.t. busbars). We use the approach by Subramanian et al. [39] to filter the space of configuration and set-actions. This produces the action space. The procedure is repeated for the different networks to generate the corresponding action spaces (see Subsection 4.1).
3.4 Regimes
We consider three environmental regimes, which represent difficulty levels for grid operation. The full-network regime involves the whole network. Lines are statically disabled in the planned-outage regime. The unplanned-outage regime introduces an opponent that disables lines spontaneously. We use the opponent specified by Manczak et al. [24]. The opponent disables a randomly selected line for four hours twice a day. A cooldown period ensures that outages are separated by at least an hour.
4 Methods
4.1 Datasets
We use the data generated with the two expert agents from our previous study [15]. The first agent is the greedy agent, which decreases the line loading greedily. The second agent is the N-1 agent, which pursues topologies that are N-1 redundant. We found that the N-1 agent performed superiorly to the Greedy agent in settings with and without outages. Both agents have an activity threshold, below which do-nothing actions are taken automatically. State-action pairs with such do-nothing actions or from unsuccessful days are excluded. Do-nothing actions above this threshold are included.
We create two datasets. First, a in-distribution (ID) dataset containing networks on which the models are trained and evaluated. This dataset was used in our previous study [15]. This dataset combines the data from the greedy agent and the N-1 agent. The data from the N-1 agent on the full network is used. The data from the greedy agent on a subset of the N-1 networks is used. These are the N-1 networks with lines 0, 2, 4, 5, 6, and 12 disabled, which the greedy agent can operate well [15].
Second, an out-of-distribution (OOD) dataset, which is used to investigate generalization. This dataset comprises data from the greedy agent on the N-1 networks with lines 1 and 3 disabled.
On both datasets, we split the datapoints into 70/10/20 train/validation/test sets based on their scenario. The sizes of the partitions are listed in Table 3. In this paper, ID or OOD networks refer to the networks in either the ID or OOD datasets. ID or OOD outages refer to the outages present in either ID or OOD networks.
[t]          Train          Val          Test          ID          196,477          26,228          59,150          OOD          41,788a          5,690a          12,724 a These datapoints are used to train a model to estimate an upper bound to generalization (see Sec. 4.8).
Table 3: The number of datapoints in the datasets and train/val/test sets.
4.2 Datapoints
Each datapoint includes the features per object, the topology vector, and the expert action. The features for each object type are listed in Table 4. The object features are normalized. If a line is disabled, its endpoint features are zero-imputed, and the corresponding topology vector values are -1. Actions are converted from a set-action format into a switch-action format. Thus, actions are represented by a vector with a length of the topology vector. Each index indicates whether the corresponding object is switched between busbars (a value of 1) or not (a value of 0). This presents a multi-label binary classification task.
Table 4: Features per object type.
Generator/Load Line endpoint
Active Production/Load Active Power Flow Current flow
Reactive Production/Load Reactive Power Flow Loading ($\rho$)
Voltage Magnitude Voltage Magnitude Thermal Limit
4.3 FCNN
The fully connected neural network (FCNN) consists of an input layer, multiple hidden layers, and an output layer. The input vector is obtained by flattening the input features into a vector and concatenating the topology vector. The hidden layers use the ReLU activation function. The output layer uses the sigmoid activation function to constrain the output to the $(0,1)$ range. The output vector has the length of the topology vector, so one value $p_{u}\in(0,1)$ is predicted per node $u$.
4.4 Homogeneous GNN
Applying a graph neural network to the power grid requires a graph representation of the power grid. We represent the grid objects as nodes. We connect nodes if the corresponding objects are attached to the same busbar or are corresponding line endpoints. We call this graph representation homogeneous as it does not consider edge types. We call the associated GNN the HomGNN. Figures 2a and 2b display an example grid and its homogeneous representation.
(a) Example Grid
(b) Homogeneous Graph Representation
(c) Heterogeneous Graph Representation
(d) Homogeneous Message Passing
(e) Heterogeneous Message Passing
Figure 2: a: An example grid consisting of two substations and seven objects. Note that this grid is unrealistic, as real power grids should not be split. b: The homogeneous graph representation of that grid. c: The heterogeneous graph representation of that grid, where edge colors indicate edge types. d: Message passing towards node 5 in the homogeneous model. The purple reflective edge is added to reflect self-weights. e: Message passing towards node 5 in the heterogeneous model, where edge colors indicate message types.
Different types of grid objects have different features (see Table 4). We use multiple two-layer perceptrons to embed the varying object features into a common embedding:
$\textbf{h}_{u,0}=\begin{cases}\text{MLP}_{gen}(\textbf{x}_{u})&\text{node $u$ represents a generator}\\ \text{MLP}_{load}(\textbf{x}_{u})&\text{node $u$ represents a load}\\ \text{MLP}_{line}(\textbf{x}_{u})&\text{node $u$ represents a line endpoint},\end{cases}$
where $\textbf{h}_{u,0}$ and $\textbf{x}_{u}$ are, respectively, the initial node embedding and the features of node $u$. The embeddings of subsequent layers are computed with the message passing rule [36]:
$\textbf{h}_{u,k+1}=\sigma(W_{self,k}\textbf{h}_{u,k}+W_{neighbor,k}\sum_{v\in\mathcal{N}(u)}\textbf{h}_{v,k}+\textbf{b}_{k}),$
where $\textbf{h}_{u,k}$ is the embedding for node $u$ in layer $k$, $\sigma$ is the activation function, $W_{self,k}$ and $W_{neighbor,k}$ are the self and neighbor weights in layer $k$, $\mathcal{N}(u)$ is the neighborhood of node $u$, and $\textbf{b}_{k}$ is the bias term. Figure 2d displays an example of homogeneous message passing. The activation function $\sigma$ is the ReLU function in all but the final layer. The final layer uses the sigmoid activation function and outputs one value. Thus, the model predicts one value $p_{u}\in(0,1)$ per node $u$.
4.5 Busbar Information Asymmetry
The aforementioned graph neural network formulation has a problem. Each node’s output specifies whether to switch the corresponding object between busbars. However, messages from objects on the current busbar are passed, while messages from the objects on the other busbar are not. We call this the busbar information asymmetry. This is intuitively problematic as a switching decision requires information about the other busbar. This also decreases the model’s expressiveness. Objects at the same substation but on different busbars and objects at entirely different substations cannot be distinguished. Moreover, the absence of inter-busbar connections also decreases the graph’s connectivity. This can result in longer paths that may require deeper GNNs.
4.6 Heterogeneous GNN
Our solution is to represent the grid with a heterogeneous graph representation. The associated GNN is called the HetGNN. This heterogeneous representation features three edge types: one for objects on the same busbar, one for objects on the other busbar (but at the same substation), and one for corresponding line endpoints. This is displayed in Figure 2c. The message passing rule is adapted to consider a weight and neighborhood aggregation per edge type:
$\textbf{h}_{u,k+1}=\sigma($ $W_{self,k}\textbf{h}_{u,k}+W_{same,k}\sum_{v\in\mathcal{N}_{same}(u)}\textbf{h}_{v,k}+$
$W_{other,k}\sum_{v\in\mathcal{N}_{other}(u)}\textbf{h}_{v,k}+W_{line,k}\textbf{h}_{line,k}+\textbf{b}_{k}),$
where $\mathcal{N}_{same}$ and $\mathcal{N}_{other}$ denote the neighborhoods of objects on the same and other busbar, respectively. The $W_{line,k}\textbf{h}_{line,k}$ term represents the addition of the embedding of a node connected by a power line. It is therefore omitted for nodes that do not represent line endpoints. Figure 2e displays an example of heterogeneous message passing.
4.7 Optimization & Postprocessing
The weights are initialized using a normal distribution, with the standard deviation set as a tuned hyperparameter. We use the Adam optimizer to minimize a label-weighted binary cross-entropy loss, defined as:
$L=$ $\textrm{mean}(-\textbf{w}(\textbf{y}\log(\textbf{p})+(\textbf{1}-\textbf{y})\log(\textbf{1}-\textbf{p}))),$
[…]
$_{i}=\begin{cases}1&\text{object $i$ corresponds to the target substation}\\ 1&\text{object $i$ corresponds to the predicted substation}\\ \alpha&\text{otherwise}\end{cases}\quad\forall w_{i}\in\textbf{w}.$ w .
There is no target substation if the target action is a do-nothing action. The predicted action is classified as a do-nothing action, without a predicted substation, if all predictions $_{i}\in\textbf{p}$ p do not exceed $0.5$. Otherwise, the predicted substation is the substation where the predictions $\textbf{p}_{s}$ at that substation $s\in S$ maximize
$\Sigma_{p_{i}\in\textbf{p}_{s}}\textrm{max}(p_{i}-0.5,0).$
Each value in the prediction vector represents whether to switch the corresponding object. However, not every prediction vector corresponds to an action in the filtered action space (see Sec. 3.3). We apply a postprocessing step that replaces the model’s prediction p with the nearest action. This postprocessing step is applied during validation, testing, and inference but not during training.
4.8 Hyperparameter Tuning & Training
The hyperparameters were tuned with two iterations of hyperparameter sweeps. This procedure was repeated for the three model types. The first sweep narrowed the hyperparameter ranges. It covered wide parameter ranges and trained for a few epochs with strict early stopping. The second sweep covered narrower ranges and trained with more epochs and less strict early stopping. All sweeps use random search with Hyperband early termination (distinct from early stopping). Hyperband early termination terminated unpromising runs early. Table 5 describes all final hyperparameter values, the second sweep ranges, and additional clarification where necessary.
Five models were trained per model type, each with different weight initializations. Each training run lasted for 100 epochs unless stopped early. Runs were terminated early if the highest validation accuracy did not increase in 20 evaluations. The validation accuracy was calculated every 50,000 iterations. The training curves are displayed in Figure 3. Finally, we trained five heterogeneous GNNs on the OOD data to contrast the models’ generalization performance. These are referred to as OOD-GNNs.
Table 5: Hyperparameters and their final sweep ranges for both the FCNNs and GNNs.
Hyperparameter
FCNN
Value
FCNN
Range
GNN
Valuea
GNN
Rangea
 #Hidden Layers 4 (1, 5) 8 (2, 12)
#Hidden Nodes 230 (32, 256) 180 (32, 256)
Batch Size 64 (32, 128) 64 (32, 128)
Learning Rate 7E-4 ($e^{-9},e^{-5})$ 2E-4 ($e^{-9},e^{-5})$
Weight Decay 0b $(e^{-9},e^{-2})$ 0b $(e^{-9},e^{-2})$
Weight Init. $\sigma$ 5 $(e^{-0.5},e^{1.5})$ 5 $(e^{-0.5},e^{1.5})$
Label Weight $\alpha$ 0.1 -c 0.1 -c
ReLU neg. slope 0.1 -d 0.1 -d
a Hyperparameter tuning was performed independently for both the HomGNN and HetGNN. However, the results were sufficiently similar. Hence, the same ranges and final values were selected. b Weight decay was set to zero after observing that the best runs had values near the lower limit. Subsequent runs with a weight decay of zero performed better. c We experimented with label weight $\alpha$ values when introducing label weighting. Label weights were not included in the hyperparameter sweeps. d The ReLU negative slope parameter was not tuned.
Figure 3: The training curves of the five models per model type. Green lines show the HetGNNs, yellow lines the FCNNs, and blue lines the HomGNNs.
4.9 Imitation Learning Agents
The ML models are applied to the environment. They can be applied directly or combined with simulation. The naive agent executes a model’s predicted action directly. We observe that this agent occasionally fails by predicting a singular erroneous action. The verify agent addresses this by verifying predicted actions with simulation. The predicted action is simulated. A do-nothing action is selected if the simulated line loadings are increased beyond the thermal limit. We also consider hybrid agents. The verify+greedy agent normally functions as the verify agent, but switches to the greedy agent if a line breaches the thermal limit. The verify+N-1 agent functions similarly but with the N-1 agent. Each agent also uses the activity threshold used by the expert agents [15].
5 Results
5.1 Supervised Learning
Table 6: The accuracies of the different models on the different combinations of data partitions and ID/OOD network sets. Numbers indicate the mean and standard deviation of accuracy for the five models. Remember that the data partitions are split by scenario.
# (%) FCNN HomGNN HetGNN OOD-GNN
 1 ID Train 79.2±0.9 75.2±0.9 83.2±1.4 -
2 ID Val 78.6±0.5 75.1±0.8 80.2±0.6 -
3 ID Test 76.6±0.6 73.0±0.6 78.5±0.8 -
 4 Raw ID Test a 73.9±0.6 67.9±0.6 74.7±0.5 -
 5 Default ID Test b 94.5±0.9 94.2±0.4 95.3±0.3 -
6 Split ID Test c 61.3±0.9 54.8±1.2 64.1±1.2 -
 7 OOD Train 35.0±0.9 63.4±1.1 67.9±0.7 86.0±0.6
8 OOD Val 37.3±0.4 65.6±0.6 69.3±0.7 83.7±0.2
9 OOD Test 34.7±0.7 61.1±0.7 65.1±0.7 80.7±0.4
a This lists the accuracy without the postprocessing step described in Sec. 4.7. b These datapoints represent the default topology, i.e., without split substations. c These datapoints represent non-default topologies, i.e., with split substations.
Table 6 lists the accuracies of the different models on the various data partitions and the ID/OOD network groups. Figure 4 shows the test accuracies of the models on both datasets. On the ID dataset (rows 1 to 3), the accuracies remain limited to approximately 80%. On each split, the HetGNN achieves the highest accuracy, followed by the FCNN and, lastly, the HomGNN. Each model type shows a tiny drop in accuracy between the train, validation, and test sets, indicating slight overfitting. Row 4 displays that accuracy drops slightly without the postprocessing step described in Sec. 4.7.
The default topology appears frequently in the ID dataset, comprising 46% of the test set. In this topology, all objects are attached to the same busbar. This avoids the busbar asymmetry problem and consequently negates the theoretical advantage of the HetGNN over the HomGNN. As shown in row 5, the accuracies on the standard topology are relatively similar. As shown in row 6, the accuracies vary considerably between topologies with split busbars.
Figure 4: The test accuracies of the model types on the ID dataset set (left) and OOD dataset (right). The ranges indicate the maximum and minimum accuracy.
The GNNs obtain far higher accuracies on the OOD dataset than the FCNNs (rows 7 to 9). The HetGNNs also obtain higher accuracies than the HomGNNs. However, the performance of either the HomGNN or the HetGNN on the OOD dataset is still substantially lower than models trained on OOD data (column ’OOD-GNN’).
5.2 Error Analysis
We repeated the error analysis previously performed on the FCNN [15] to the two types of GNN models. We again find that the class imbalance and class overlap play a role in the limited accuracy. Figure 5 shows that both GNN types predict infrequent classes disproportionately infrequently. We investigated pairs of classes that were often confused, i.e., pairs of classes that were often mistaken for one another in prediction. Figure 6 shows that the errors of frequently confused classes are in overlapping regions.
Figure 5: The log-distributions of the classes in the ID validation set. Overlaid is the frequency by which the model predicts that class (left: HomGNN, right: HetGNN). The non-overlapping blue areas at the tails of the distributions indicate that the models predict rare classes disproportionally infrequently. This finding is consistent among the models.
Figure 6: The datapoints of two classes that are most often confused, for the N-1 network with line 0 disabled. The datapoints are projected on the first two principal components. The datapoints confused by the models (left: HomGNN, right: HetGNN) are overlaid in red. As visible, the datapoints in the overlapping region are confused. This finding is consistent among the models and confused classes.
Inspection of the nearest neighbors supports the notion that the class overlap contributes to low accuracy. We applied one of the HetGNNs to a subset of 2,500 validation data points from the full network. The accuracy over the data points whose nearest neighbor is in the same class is 92.97%, but only 44.44% for those whose nearest neighbor is in another class. Other models showed similar results.
5.3 Graph Smoothness
The tendency of graph neural networks to suffer from oversmoothing is well-documented [3, 23]. We compute the mean average distance (MAD) values to measure graph smoothness [3]. MAD values quantify the similarity of node embeddings. Small MAD values indicate that node embeddings are similar and, thus, that the graph is smooth.
The MAD values are computed between neighbor nodes over the first 1000 validation datapoints. Figure 5.4 shows the MAD values between the various layers. As visible, the MAD values in the middle GNN layers are consistently higher in the HetGNN than in the HomGNN. This suggests that the heterogeneous representation leads to a less smooth local embedding. This finding is consistent among the models.
5.4 Graph Diameter
We mentioned that the heterogeneous graph representation can lead to shorter paths. We measure this by the diameter, i.e., the shortest path length between the two most distant nodes. We calculate the diameter by finding the lowest exponent of the adjacency matrix that produces a matrix without zero entries [7]. The diameters of the 25 most common topologies are shown in Figure 5.4. As visible, the heterogeneous graph representation infrequently leads to a shorter diameter.
[…]
figureThe average neighbor MAD values per layer of a HomGNN and a HetGNN. As visible, the HomGNN has far lower neighbor MAD values in the middle layers. Layers 0 and 8 indicate the neighbor MAD values before and after the message-passing layers. Error bars indicate the standard deviations over the 1000 datapoints.
[…]
figureThe diameter of the 25 most common topologies. The diameter of the heterogeneous graph representation is invariant to busbar reconfiguration.
5.5 Simulation Performance
[…]
Table 7: The percentage of days completed by agents in various regimes. The agents were evaluated on the test scenarios.
Table 7 shows the days completed by the various agents in the different settings. The ’Full Network’ column describes the setting without outages. The ’ID Outages’ describes a setting with an opponent that causes random outages present in the ID dataset (see Sec. 4.1). The ’OOD Outages’ describes a setting with an opponent that causes random outages present in the OOD dataset.
The agents can be ordered by performance: Do-Nothing $<$ Naive $<$ Verify $<$ Greedy Hybrid $<$ N-1 Hybrid. The naive agents (rows 4 to 6) and verify agents (rows 8 to 10) come close to the performance of the N-1 expert agents (row 3). The hybrid agents (rows 12 to 14, and 16 to 18) can match the N-1 expert agents.
On the setting with no or ID outages (columns ’Full Network’ and ’ID Outages’) the HetGNN network consistently performs best, followed by the FCNN and, finally, the HomGNN (rows 4 to 6, and 8 to 10). Additional simulation reduces or nullifies the differences (rows 12 to 14, and 16 to 18).
On the setting with OOD outages (column ’OOD Outages’), the GNNs outperform the FCNN (rows 4 to 6, and 8 to 10). However, additional simulation also decreases this effect (rows 12 to 14, and 16 to 18). Furthermore, the performance of the models trained on the ID dataset remains substantially lower than models trained on OOD dataset (rows 7 and 11), even with additional simulation (rows 15 and 19).
5.6 Efficiency
Finally, we consider the inference speed of the different agents. The inference speed was measured on an Apple M1 CPU with minimal background processes on the first fifty validation scenarios. Table 5.6 lists the durations per model type and agent. Figure 5.6 plots these values against the agent’s performance on the full-network topology. The GNNs, particularly the HetGNNs, are considerably slower than the FCNNs. However, these differences are small relative to the simulation times of the expert agents, which are orders of magnitude higher. Most importantly, all GNN agents have a favorable combination of speed and performance.
[…]
figureThe performance and inference duration for the agents on the setting with OOD outages. The combinations of speed and performance of the imitation learning agents are favorable. All GNN agents are on the Pareto front (dashed line). The GNN models are somewhat slower than the FCNN models.
[…]
tableInference durations of various agents. The hybrid models all use the FCNN. The differences between model types for the hybrid agents are relatively insignificant. Agent Duration ($\upmu$s) Greedy 6.77E5 N-1 4.70E6 Naive FCNN 5.74E2 Naive HomGNN 2.81E3 Naive HetGNN 3.84E3 Verify FCNN 1.14E4 Verify HomGNN 1.50E4 Verify HetGNN 1.58E4
6 Discussion
The heterogeneous GNN consistently performs best, followed by the FCNN, and lastly, the homogeneous GNN. This is true for both the accuracy and grid operation. The superiority of the HetGNN to the FCNN indicates that GNNs can exploit the graph structure better. The inferiority of the HomGNN can be attributed to the busbar asymmetry problem. However, smoother graphs and longer path lengths might also play a role. The HetGNN’s coarser graphs can be attributed to the more diverse edge representation. The GNNs take considerably longer to evaluate than the FCNNs, although this difference is very tiny compared to the expert agents.
The GNNs achieve far higher accuracy than the FCNNs on the OOD networks. This is also reflected in the simulation performance, although the effect is smaller. Although GNNs show a superior ability to generalize to OOD networks, they still perform considerably worse than GNNs trained on these networks.
There seemed to be a limit to the accuracy of the trained models. This can be attributed to class imbalance and class overlap. We hypothesize that the class overlap originates from the Grid2Op simulate function. It is possible that identical network states have diverging forecasts and consequently diverging actions. The limited accuracy does not stop the ML models from successfully completing many of the grid operation days. The naive ML models come close to the performance of the expert agents. The hybrid agents can match them.
Hybrid agents can outperform expert agents in specific settings, which is surprising. We hypothesize that this results from omitting the state-action pairs from failed runs. The model bias towards frequent actions might also result in more robust topologies.
An interesting effect is that model differences are more pronounced in accuracy than simulation performance. Similarly, the limited accuracy does not stop the models from operating the grid well. We have two explanations for this. Firstly, a misprediction does not necessarily lead to a game-over. This is probably particularly true in regions with class overlap. Secondly, difficult days led to the generation of many relatively uncommon datapoints. Failure to predict many such datapoints would consequently affect the completion of only few days.
7 Future Work
Our findings indicate the promise of fully graphical neural networks for topology control methods. However, further research is necessary to establish this. Future work should compare fully GNN models with models that only use GNNs for feature extraction. Similarly, the heterogeneous graph representation should be introduced into other approaches. Moreso, it would be good to apply these methods in a reinforcement learning framework.
Further efforts are likely also necessary to scale this approach to larger grids. Fully GNN models might require very deep GNNs to model large graphs. Very deep GNNs are associated with additional challenges [46]. Future research could focus on methods to downsample grid graph representations, e.g., through trainable graph pooling [12]. Alternatively, it might be beneficial to combine GNN layers and other layers to capture long-range patterns.
The imitation learning approach used here can also be developed further. Our approach suffers from distribution shift, a compounding divergence between the behavior of the ML and the expert agent [1]. More advanced IL frameworks, such as DAgger, address this [34]. Investigating other expert agents and developing an improved dataset would also be valuable.
Finally, the generalization capabilities of graph neural networks should be studied more extensively. Future studies should investigate how factors such as the diversity of training data and differences between the ID and OOD data affect generalization.
8 Statements
8.1 CRediT authorship contribution statement
Matthijs de Jong: Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Visualization. Jan Viebahn: Conceptualization, Methodology, Resources, Writing - Review & Editing, Supervision, Project Administration. Yuliya Shapovalova: Methodology, Resources, Writing - Review & Editing, Supervision.
8.2 Declaration of competing interest
Matthijs de Jong reports a relationship with TenneT TSO BV that includes: employment. Jan Viebahn reports a relationship with TenneT TSO BV that includes: employment. If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8.3 Acknowledgements
Special thanks to Mohamed Hassouna from Fraunhofer Society for providing valuable feedback. This work is supported by Graph Neural Networks for Grid Control (GNN4GC) funded by the Federal Ministry for Economic Affairs and Climate Action Germany under the funding code 03EI6117A.
8.4 Funding
AI4REALNET has received funding from European Union’s Horizon Europe Research and Innovation programme under the Grant Agreement No 101119527. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
8.5 Code & Data Availability
Our code \@footnotemark and data [16] are publicly available.

### v3

@@
-Generalizable Graph Neural Networks for Robust Power Grid Topology Control
+Graph Neural Networks for Transmission Grid Topology Control: Busbar Information Asymmetry and Heterogeneous Representations
 Abstract
-The energy transition necessitates new congestion management methods.
-One such method is controlling the grid topology with machine learning (ML).
-This approach has gained popularity following the Learning to Run a Power Network (L2RPN) competitions.
-Graph neural networks (GNNs) are a class of ML models that reflect graph structure in their computation, which makes them suitable for power grid modeling.
-Various GNN approaches for topology control have thus been proposed.
-We propose the first GNN model for grid topology control that uses only GNN layers.
-Additionally, we identify the busbar information asymmetry problem that the popular homogeneous graph representation suffers from, and propose a heterogeneous graph representation to resolve it.
-We train both homogeneous and heterogeneous GNNs and fully connected neural networks (FCNN) baselines on an imitation learning task.
-We evaluate the models according to their classification accuracy and grid operation ability.
-We find that the heterogeneous GNNs perform best on in-distribution networks, followed by the FCNNs, and lastly, the homogeneous GNNs.
-We also find that both GNN types generalize better to out-of-distribution networks than FCNNs.
+Factors such as the proliferation of renewable energy and electrification contribute to grid congestion as a pressing problem.
+Topology control is an appealing method for relieving congestion, but traditional approaches for topology discovery have proven too slow for practical application.
+Recent research has focused on machine learning (ML) as an efficient alternative.
+Graph neural networks (GNNs) are particularly well-suited for topology control applications due to their ability to model the graph structure of power grids.
+This study investigates the effect of the graph representation on GNN effectiveness for topology control.
+We identify the busbar information asymmetry problem inherent to the popular homogeneous graph representation.
+We propose a heterogeneous graph representation that resolves this problem.
+We apply GNNs with both representations and a fully connected neural network (FCNN) baseline on an imitation learning task.
+The models are evaluated by classification accuracy and grid operation ability.
+We find that heterogeneous GNNs perform best on in-distribution network configurations, followed by FCNNs, and lastly, homogeneous GNNs.
+We also find that both GNN types generalize better to out-of-distribution network configurations than FCNNs.
 keywords:
@@
 {highlights}
-Heterogeneous over homogeneous graphs increase graph neural network expressiveness
-Heterogeneous graph neural networks are the highest-performing models
-Graph neural networks generalize better to out-of-distribution networks
-Machine learning agents obtain good performance at high efficiency
+Busbar information asymmetry limits graph neural networks for grid topology control.
+Heterogeneous graph models of the grid increase model expressiveness and performance.
+Machine learning agents obtain good grid operation performance at high efficiency.
+Graph neural networks generalize better to out-of-distribution grid states.
 1 Introduction
-The energy transition is crucial for ensuring society’s sustainability and future security.
-Transmission system operators (TSOs) play a crucial role in this transition.
-They are consequently facing new operational challenges [9].
-Grid congestion is a problem that TSOs are already encountering but which the energy transition will exacerbate [42].
-If not mitigated, grid congestion could make the future availability of energy unreliable.
-There are several actions that TSOs can take to alleviate grid congestion [17].
-Presently, redispatching and curtailment are the primary methods.
-However, these actions are often costly, disruptive, or not applicable.
-Topology control presents an opportunity for remedying grid congestion without these problems [41].
-Congestion can be relieved by exploiting the flexibility in the network topology.
-However, this technique mostly remained unutilized, as the space of potential topologies is combinatorially large.
-Due to the recent congestion management challenges, there has been a new interest in topology control.
-This interest led to the development of the GridOptions tool for remedial action recommendation by Viebahn et al.
-[41].
-This tool validates the practical value of topology control.
-Even simple topological strategies could reduce line congestion by ten to twenty percent.
-The recommended strategies improved upon strategies without topological actions or those considered by grid operators.
-However, while the results are promising, further developments are necessary for broader applicability.
-The tool was only applied to a small subset of the Dutch power grid, featuring nine substations.
-Additionally, the inference time of the tool is relatively long: one hour.
-The next challenges involve scaling this approach and reducing inference time.
-Machine learning (ML) presents a solution to these challenges [42].
-ML models can learn which topologies are suitable in which situations and predict these topologies much quicker than traditional computational methods.
-To facilitate research in this area, the Learning to Run a Power Network (L2RPN) competitions were hosted [28, 27].
-Furthermore, the Grid2Op Python library was developed.
-[^1] https://grid2op.readthedocs.io/ This library greatly simplified the development and evaluation of ML methods for grid topology control.
-These developments kick-started research into ML methods for grid topology control.
-Classical ML models are typically designed for tabular data and hence fail to capture the graph structure of power grids.
-This is problematic as the grid topology frequently changes, e.g., because of maintainance.
-Power grid operation requires adapting to such changes.
-Graph neural networks (GNNs) [36] are a class of ML models for graph-structured data.
-The computation in GNNs can reflect arbitrary graphs, which allows them to exploit a graph’s structure.
-This makes GNNs suitable for modeling power grids.
-GNNs have thus been applied to many grid operation tasks [5, 4, 31, 11, 18, 13, 30].
-Various GNN approaches have been proposed for grid topology control [45, 43, 44, 35, 40, 33].
-This research has demonstrated the feasibility of GNN approaches for grid topology control.
-However, important topics have not yet been addressed:
+The energy transition, while critical, introduces significant new challenges to transmission system operators (TSOs) entseo_energy_transistion .
+The inherent variability and limited dispatchability of renewable energy sources can lead to supply peaks.
+TenneT, the Dutch TSO, projects an increase in installed renewable power capacity in the Netherlands from 29.6 GW in 2022 to 103.8 GW by 2033 tennet_2024 .
+Simultaneously, the electrification of sectors such as transportation is driving an increase in total electrical demand.
+TenneT forecasts Dutch electrical demand to increase from 113 TWh in 2022 to 171 TWh in 3033, representing a growth of over 50% tennet_2024 .
+Similar trends are anticipated in neighboring countries tennet_2024 .
+These developments introduce challenges such as grid congestion.
+Grid congestion occurs when grid components approach or exceed their operational capacity limits.
+Congestion already poses a problem with tangible consequences: the European Union Agency for the Cooperation of Energy Regulators estimated the cost of European grid congestion in 2023 to be approximately 4 billion EUR ACER2024_CrossZonalCapacities .
+Furthermore, potential grid congestion scenarios can lead to the denial of connection requests for prospective customers, halting the development of businesses, schools, housing, and other facilities.
+These problems will quickly escalate without intervention.
+Beyond the development of new infrastructure and the rejection of connection requests, TSOs can employ other measures.
+TSOs typically use redispatching and/or curtailment, involving the rebalancing or reduction of power production.
+However, both approaches are costly and limited in their applicability.
+Topology control presents an alternative method for congestion management that is cost-effective and flexible.
+Network topology can be altered by reconfiguring component connections within substations and by disabling power lines.
+The resulting change in power flow can alleviate congestion.
+Topology control thus present a cheap and flexible congestion management technique kelly2020rlelec .
+Nevertheless, the usage of topology control by TSOs remains minimal.
+The number of possible network topologies grows exponentially with grid size, and most topologies cannot be operated.
+Moreover, topological changes affect power flow in complex, nonlinear ways that are time-consuming to compute.
+Finding the correct topology for a particular situation is thus difficult, and usage of topology control has hence remained minimal.
+However recent advancements have been made with the introduction of GridOptions by TenneT, a pioneering real-world tool for topological strategy suggestion jan_viebahn_2024 .
+The tool has demonstrated the value of topology control, showing that even simple topological strategies could reduce line congestion by 10 - 20%.
+The tool currently relies on fast load-flow calculations for screening topologies.
+However, the tool’s scope remains limited to a small section of the Dutch grid, including only eight substations, and the long inference time limits it to day-ahead application.
+To remove these limitations, the developers propose the future use of machine learning to speed up computation.
+The idea of mitigating grid congestion with topology control dates back to the 1980s 4334990 .
+The discrete, nonlinear nature of topological actions limits the applicability of direct optimization methods and early research therefore often relied on computationally intensive simulations of topological actions.
+To address time constraints, simplified models or solvers such as linearized DC approximations were frequently used 515208 , 6939408 .
+Advanced methods have used more sophisticated solvers 1525118 or techniques to narrow the set of topological actions considered 7792703 .
+Nonetheless, these approaches often rely on inaccurate approximations, and their reliance on simulation renders them infeasible for large-scale real-world usage.
+Other research has focused on mathematical optimization techniques, such as mixed-integer programming (MIP) to approximate optimal busbar configurations.
+However, such approaches also prove too slow, with proposed approaches being limited to only line-switching 4492805 or busbar-switching within a single substation 9483070 .
+Recently, machine learning (ML) techniques have garnered significant interest within the power systems community.
+ML models are notoriously good at learning complex nonlinear patterns, such as the patterns caused by topological actions.
+Furthermore, inference with ML models is quick as it does not require extensive search or optimization.
+The Learning to Run a Power Network (L2RPN) competitions marot2019learning , marot2021learning and the Grid2Op Python library grid2op  [^1] https://grid2op.readthedocs.io/ played a large part in the popularization of this research area.
+The field now has an active research community.
+Power grid operation is ultimately a sequential decision-making problem.
+ML research on topology control has largely focused on the reinforcement learning Subramanian_2021 , 10.1609/aaai.v37i12.26724 and imitation learning lan2019aibased , lehna2023managing paradigms.
+Reinforcercement learning (RL) involves training agents by directly interacting with the environment.
+However, RL models for topology control often converge to “do-nothing” policies because the default topology is mostly satisfactory and topologies that improve upon it are rare githubGitHubAsprinChinaL2RPN_NIPS_2020_a_PPO_Solution .
+Imitation learning (IL) involves training a ML model to replicate the behavior of an expert agent.
+It can be used either as pre-training for RL to avoid the aforementioned problem or as a stand alone method deJong2024 .
+Evolutionary strategies have been also been proposed for the problem zhou2021actionsetbasedpolicy .
+Graph neural networks (GNNs) 4700287 have recently recieved much attention, both from the machine learning and power engineering community.
+Deep learning models have traditionally been designed for tabular, sequential, or grid-like data.
+Consequently, ML methods for power grid operation have historically not or minimally exploited the network structure of power grids.
+This is a major limitation as power grid topologies are dynamic and frequently change, for example due to maintenance.
+Graph neural networks encode network structure into their representation and are therefore well-suited for power grid modeling.
+GNNs have hence been applied to various tasks related to grid operation, such as solving power flows 8851855 , DONON2020106547 , optimal power flow 9053140 , grid reliability or risk assessment CAMBIERVANNOOTEN2025125401 , ZHANG2025124793 , and state estimation NGO2024122602 .
+Similarly, GNNs have been applied to the problem of topology control with promising results yoon2020winning , 9830198 .
+However, despite the existing research, important topics have not yet been addressed by the literature:
   1. 1.
-    All approaches that directly predict topology actions with GNNs only use GNN layers as feature extractors or embedders [45, 43, 44, 33].
-Fully-connected or attention layers follow the GNN layers.
-No research has investigated ’fully graphical’ neural networks for grid topology control.
-Purely GNN approaches might be able to exploit the topology more deeply.
+    Most present literature applying graph neural networks to topology control uses the same graph representation 9535409 , vandersar2023 , 9347305 , 9830198 , Hassouna2025LearningTA .
+In this representation, configurable objects are connected by edges if they are connected to the same busbar or power line.
+This is, however, only one of many ways to represent the power grid as a graph.
+On the related problem of optimal power flow, progress has been made by considering richer, heterogeneous graph representations ghamizi2024powerflowmultinet .
+However, for topology control, a study comparing or proposing improved graph representations for GNNs does not yet exist.
   2. 2.
-    All present literature uses the graph representation specified by which busbars and power lines objects share.
-However, this is only one of many ways to represent the power grid as a graph.
-There has not been any study that compares graph representations and their effect on GNN performance.
-  3. 3.
-    Grid changes are frequent, and operation needs to continue despite these changes.
-For robust operation, models must generalize to potentially unexpected network states.
-A previous competition has focused on generalization to out-of-distribution injection profiles [27].
-However, no work has investigated generalization to out-of-distribution networks.
-This study aims to fill the aforementioned gaps in applying GNNs to grid topology control. We develop fully GNN models that use a richer graph representation. The main contributions of this study are:
+    Power grids are frequently subject to changes, which cannot impede operation.
+Reliable and robust operation requires models to generalize or adapt to, potentially unexpected, altered network states.
+Prior work has focused on generalization to out-of-distribution injection profiles marot2021learning .
+However, no work has investigated generalization to out-of-distribution network configurations.
+This study addresses these gaps in the literature.
+We identify a problem with the popular graph representation and propose a heterogeneous graph representation that resolves it.
+We train the GNNs, as well as ‘traditional’ fully connected neural networks (FCNNs), on a multi-label binary classification imitation learning task.
+The utilized dataset consists of various topologies and outages deJong2024 .
+The models are evaluated by their classification accuracy and their efficacy when used as agents to operate a simulated power grid.
+Furthermore, we compare the operation efficacy and inference speed of ML models against expert and hybrid agents.
+We also investigate the accuracy and operation efficacy of models to out-of-distribution network configurations.
+In particular, the foremost contributions of this study are:
   1. 1.
-    We propose the first fully graphical neural network approach for grid topology control.
-The models are trained on a multi-label binary classification imitation learning task involving various networks and topologies.
-We compare the GNNs against fully connected neural networks (FCNNs).
-We evaluate the models by their classification accuracy and ability to operate the power grid as agents.
-We also consider expert and hybrid agents and compare their inference speed.
+    We perform the first analysis of the effect of power grid graph representation on GNN topology control performance.
+The currently popular graph representation models only the current connections between objects.
+However, it ignores which objects might be connected by actions, thereby obscuring the effects of actions.
+We call this the ‘busbar information asymmetry’, which hinders information flow and limits GNN expressiveness.
+We propose a heterogeneous graph representation that models both current and potential connections and consequently resolves the asymmetry.
+We evaluate both graph representations with GNNs.
   2. 2.
-    We first analyze the effect of grid graph representation on GNN performance.
-The graph representation in the literature models current but not potential object-busbar attachments.
-This results in busbar information asymmetry, which blocks information flow and impairs GNN expressiveness.
-We propose a heterogeneous graph representation that models current and potential connections, which remedies busbar information asymmetry.
-We evaluate both graph representations with GNNs.
-  3. 3.
-    We first investigate the ability of graph neural networks to generalize topology control to out-of-distribution grids.
-Generalization is evaluated w.r.t.
-model accuracy and the ability of agents to operate the power grid, and contrasted against FCNNs.
+    We perform the first investigation of the ability of graph neural networks to generalize topology control to out-of-distribution grid configurations.
+We evaluate this both by model accuracy and grid operation effectiveness.
 This paper is organized as follows.
-Section 2 describes related work in ML grid topology control.
 Section 2 describes the power grid setup used in this study.
@@
 Section 4 describes the evaluation and analysis of the ML models.
-Sections 6 and 7 provide discussion and recommendations for future work, respectively.
-The code for this project is available.
+Sections 5 provides discussion and recommendations for future work, respectively.
+The code for this project is available online.
 [^2] https://github.com/MatthijsdeJ/GNN_PN_Imitation_Learning Table 1 lists the meaning of mathematical symbols used in this paper.
-Table 2 lists the glossary.
 Table 1: Mathematical symbols used in this paper and their meaning.
 Notation Meaning
-$O$ The set of grid objects (i.e., generators, loads, and line endpoints).
-$S$ The set of substations.
-$\textbf{x}_{u}$ The features of a GNN model at node $u\in O$.
-$\textbf{h}_{u,k}$ The node embeddings of a GNN model at node $u\in O$ in layer $k$.
-$\mathcal{N}(u)$ The neighbors of node $u\in O$.
-$W$ A weight matrix.
-$\sigma$ An activation function.
-$\textbf{p}=(0,1)^{|O|}$ The output of a ML model.
-$\textbf{y}=\{0,1\}^{|O|}$ The target of a ML model.
-$\alpha=0.1$ The label weight hyperparameter.
-$\textbf{w}=\{\alpha,1\}^{|O|}$ The label weights for the loss.
-$\{\textbf{p}_{s}\}{}\forall s\in S$ A partition of the output vector p into subvectors $\textbf{p}_{s}$. Subvector $\textbf{p}_{s}$ corresponds to substation $s\in S$.
-$\eta$ = 0.97 The activity threshold parameter.
-$\theta=1.0$ The risk threshold parameter, used by various agents.
-Table 2: Glossary
-Term Meaning
-Busbar Information Asymmetry An asymmetry in the availability of information of objects on busbars.
-Full-network Regime An environmental regime without statically or randomly disabled lines.
-In-distribution Data (ID) Data with networks that the models are exposed to during training.
-Injection The generators and loads of the environment.
-Loading The ratio of a line’s current over its thermal limit.
-Multi-label Binary Classification A learning task where the target consists of multiple binary values.
-N-1 Network A variation of the default network with a single line disabled.
-N-1 Redundancy The state of the network being resilient to the disablement of any single line.
-Object Generators, loads, and line endpoints in the Grid2Op environment.
-Out-of-distribution Data (OOD) Data with networks that the models are not exposed to during training; used to investigate generalization.
-Planned-outage Regime An environmental regime with statically disabled lines.
-Topology Reversal The periodical resetting of the topology to the default topology.
-Topology Vector The Grid2Op variable that specifies the object-busbar configurations.
-Unplanned-outage Regime An environmental regime with an opponent that randomly disables lines.
-2 Related Work
-The L2RPN competitions piqued the academic interest in grid topology control with machine learning.
-Prior research into topology control used linear programming [10, 49], but such methods are too rigid and/or slow to be practical.
-The first L2RPN competition premiered in 2019.
-It focused on operating scenarios of the EEE14-bus network within a specific time frame [28].
-The following L2RPN competition was featured at WCCI in 2020 and featured the larger IEEE118-bus network [29].
-The next competition at NeurIPS 2020 [27] introduced a robustness track with unplanned outages and an adaptability track with out-of-distribution injection profiles.
-Later L2RPN competitions focused on trust and sending an alarm signal [26], and the inclusion of batteries [37].
-Graph neural networks have been successfully applied to many tasks of power grid operation, among which power flow estimation [5, 4], optimal power flow computation [31, 11], load shedding [18], outage prediction [13], and stability prediction [30].
-For a comprehensive survey on graph reinforcement learning for power grids, we recommend the survey by Hassouna et al.
-[14].
-Several studies investigated GNNs for topology control.
-Yoon et al.
-[45] propose a hierarchical actor-critic approach that uses a GNN module to obtain graph embeddings.
-Their approach won the L2RPN WCCI 2020 challenge.
-Xu et al.
-[43] are the first to apply GNNs with graph attention layers for grid topology control.
-They apply this GNN with a proximal policy optimization approach with an action space searched with monte-carlo tree search.
-Their approach can, on average, maintain the system for four days in the 2020 NeurIPS L2RPN robustness track.
-Xu et al.
-[44] find that a deep reinforcement learning agent which uses features extracted with a GNN trains more quickly and achieve better performance than an agent without.
-Van der Sar et al.
-[35] propose a hierarchical actor-critic approach with one agent per substations.
-These agents use graph neural networks blocks, although the details are left unspecified.
-They apply this approach to the IEEE5-bus network.
-Qiu et al.
-[33] propose a method similar to Yoon et al.’s [45] and obtain similar results.
-Taha et al.
-[40] propose a different approach than directly predicting topology actions with GNNs.
-They use GNNs for grid estimation.
-These estimates are then used in grid topology tree search.
-We will also discuss the approach by Zhao et al.
-[47], which is relevant because of their focus on generalization.
-They propose a proximal policy optimization for redispatching that relies on graph embeddings.
-The graph embeddings are generated by a GraphSAGE model trained with a unsupervised learning task.
-They find that this graph module can succesfully represent out-of-distribution grid changes.
-Other 2020 NeurIPS L2RPN submissions also used graph neural networks but were not notably successful [27].
-Other work did not use graph neural networks.
-These approaches have mostly used reinforcement learning (RL).
-Chauhan et al.
-[2] propose the PowRL framework, which integrates a novel heuristic with an RL approach.
-They achieved the best results on the legacy 2020 L2RPN robustness track yet.
-Dorfer et al.
-[6] adapt the AlphaZero framework [38] for the power grid domain.
-They achieved top results in the legacy 2020 WCCI L2RPN competition.
-Pan et al.
-[32] investigate the vulnerability of RL methods to adversarial attacks.
-They propose a more robust agent trained adversarially.
-Subramanian et al.
-[39] perform an in-depth investigation of a simple agent on a single scenario.
-Manczak et al.
-[24] propose a hierarchical RL approach that separates substation and configuration selection.
-Imitation learning (IL) has also been applied to the problem of grid topology control.
-De Jong et al.
-[15] recently performed the first investigation in IL as a stand-alone method for topology control.
-They found that both IL and hybrid agents could obtain good performance and inference speed combinations.
-Prior work on IL used IL as a ’warm start’ for RL.
-Lan et al.
-[20] first proposed IL to pre-train an RL model for grid control.
-Binbinchen [8] achieved second place in the 2020 NeurIPS L2RPN submission by expanding this approach.
-Lehna et al.
-[22, 21] refined this approach.
-They introduced topology reversal and a higher-performing N-1 expert agent [22], and a heuristic target topology approach [21].
-The winning submission of the 2023 L2RPN competition also included this approach [19].
-Other approaches used neither RL nor IL.
-Zhou et al.
-[48] use evolutionary strategies to train a model to narrow down the action space.
-The submission that won both 2020 NeurIPS L2RPN tracks used a similar approach [27].
-Although not ML, the expert system developed by Marot et al.
-has also achieved notoriety [25].
+$O$ Set of grid objects (i.e., generators, loads, and line endpoints).
+$S$ Set of substations.
+$\textbf{x}_{u}$ Features of a GNN model at node $u\in O$.
+$\textbf{h}_{u,k}$ Node embeddings of a GNN model at node $u\in O$ in layer $k$.
+$\mathcal{N}(u)$ Neighbors of node $u\in O$.
+$W$ Weight matrix.
+$\sigma$ Activation function.
+$\textbf{p}=(0,1)^{|O|}$ Output of a ML model.
+$\textbf{y}=\{0,1\}^{|O|}$ Target of a ML model.
+$\alpha=0.1$ Label weight hyperparameter.
+$\textbf{w}=\{\alpha,1\}^{|O|}$ Label weights for the loss.
+$\{\textbf{p}_{s}\}{}\forall s\in S$ Partition of the output vector p into subvectors $\textbf{p}_{s}$. Subvector $\textbf{p}_{s}$ corresponds to substation $s\in S$.
+$\eta$ = 0.97 Activity threshold parameter.
+$\theta=1.0$ Risk threshold parameter, used by various agents.
 2 Power Grid Setup
 2.1 Power Grid
-In this study, we experiment on the IEEE14-bus system, displayed in Figure 1.
-We simulate this system with the rte_case14_realistic environment from the Grid2Op library [^3] This environment has since been deprecated.
+The IEEE 14-bus system (Fig.
+1) was used for the experiment in this study.
+Simulations were conducted using rte_case14_realistic environment of the Grid2Op library [^3] Note that this environment has since been deprecated.
 Future research should use environment l2rpn_case14_sandbox..
-The power grid includes fourteen substations, five generators, 11 loads, and 20 power lines.
-One generator is solar, one nuclear, one wind-based, and two are thermal.
-The grid is divided into two sides.
-The high-voltage transmission side contains substations 0 to 4.
-The low-voltage distribution side contains substations 5 to 13.
-Lines 15 to 19 model the transformers connecting the two sides of the grid.
-We adjust the thermal limits to the values specified by Subramanian et al.
-[39].
-This makes the transmission and distribution differences more pronounced and realistic.
-We also investigate network variations with single lines disabled.
-We refer to these as N-1 networks.
-We use the N-1 networks to investigate the ability of agents to operate with line outages and to test the generalization of ML models to out-of-distribution networks.
-Figure 1: The default state of Grid2Op environment rte_case14_realistic.
+The system comprises 14 substations, five generators (one solar, one nuclear, one wind-based, and two thermal), 11 loads, and 20 power lines.
+Substations 0-4 represent the high-voltage transmission side, while substations 5-13 represent the low-voltage distribution side, interconnected by transformers modeled by lines 15-19.
+To amplify the differences between transmission and distribution characteristics, thermal limits were adjusted according to the values specified by Subramanian et al.
+Subramanian_2021 .
+Furthermore, we investigate network variations/configurations with single lines disabled, referred to as N-1 networks.
+These N-1 networks were used to investigate the ability of agents to operate with line outage conditions and to assess the generalization of ML models to out-of-distribution network configurations.
+Figure 1: The default state of Grid2Op environment rte_case14_realistic. Numbers indicate power line or substation IDs.
 2.2 Operational Period
-The environment features one thousand scenarios, each consisting of 8064 five-minute timesteps.
-One scenario thus models 28 days.
-Each scenario features different injection profiles.
-We split the scenarios into individual days.
-This daily period better reflects operational periods and emulates beneficial topology reversal [22].
-A game-over occurs when the power grid fails to transport sufficient power from generators to loads [^4] Three scenarios were excluded where Grid2Op failed to converge the power flow without agent misoperation.
+The environment features 1000 distinct scenarios, each spanning 8064 five-minute timesteps, thus modeling a 28-day period.
+Each scenario is characterized by unique injection profiles.
+We split the scenarios into individual days, creating 28,000 episodes.
+This day-long period better reflects operational periods, and resetting the topology between days emulates beneficial topology reversal lehna2023managing .
+A simulation terminates early (“game over”) when the power grid is unable to transport sufficient power from generators to loads [^4] In three scenarios, the Grid2Op power flow solver failed to converge, not as a result of agent misoperation.
 These scenarios were omitted throughout the study..
 2.3 Action Space
-The substations have precisely two busbars in the environment.
-The topology actions switch objects, i.e., generators, loads, and line endpoints, between these busbars.
-The topology vector specifies the busbar attachments of objects.
-Each index in the vector corresponds to an object.
-At an index, a value of 1 indicates attachment to the first busbar, 2 indicates attachment to the second busbar, and -1 indicates object disconnection.
-Actions are limited to a single substation per timestep to reflect real-world timing constraints.
-Certain substation configurations are invalid (isolated generators or loads) or redundant (configurations mirrored w.r.t.
-busbars).
-We use the approach by Subramanian et al.
-[39] to filter the space of configuration and set-actions.
-This produces the action space.
-The procedure is repeated for the different networks to generate the corresponding action spaces (see Subsection 4.1).
+Within the simulation environment, each substation is modeled with two busbars.
+Topology control actions involve switching network objects (generators, loads, and line endpoints) between these busbars.
+A topology vector specifies the busbar attachment of each object; each index in the vector corresponds to a specific object, with a value of 1 indicating connection to the first busbar, 2 indicating connection to the second busbar, and -1 indicating disconnection.
+Actions are limited to a single substation per timestep, reflecting real-world operational constraints.
+Certain substation configurations are considered invalid (e.g., isolated generators or loads) or redundant (configurations mirrored across busbars).
+The action space is generated by applying the filtering approach developed by Subramanian et al.
+Subramanian_2021 .
+The procedure is repeated for each network variation to generate the corresponding action spaces (see Subsection 3.1).
 2.4 Regimes
-We consider three environmental regimes, which represent difficulty levels for grid operation.
-The full-network regime involves the whole network.
-Lines are statically disabled in the planned-outage regime.
-The unplanned-outage regime introduces an opponent that disables lines spontaneously.
-We use the opponent specified by Manczak et al.
-[24].
-The opponent disables a randomly selected line for four hours twice a day.
-A cooldown period ensures that outages are separated by at least an hour.
+To evaluate performance under varying operational stress, three environmental regimes were considered.
+The full-network regime does not involve line outages.
+The planned-outage regime incorporates static line outages.
+The unplanned-outage regime simulates stochastic line outages using the opponent described by Manczak et al.
+manczak2023hierarchical .
+This opponent disables a randomly selected line for a four-hour duration, twice daily, with a minimum one-hour interval between outages.
 3 Methods
 3.1 Datasets
-We use the data generated with the two expert agents from our previous study [15].
-The first agent is the greedy agent, which decreases the line loading greedily.
-The second agent is the N-1 agent, which pursues topologies that are N-1 redundant.
-We found that the N-1 agent performed superiorly to the Greedy agent in settings with and without outages.
-Both agents have an activity threshold, below which do-nothing actions are taken automatically.
-State-action pairs with such do-nothing actions or from unsuccessful days are excluded.
-Do-nothing actions above this threshold are included.
-We create two datasets.
-First, a in-distribution (ID) dataset containing networks on which the models are trained and evaluated.
-This dataset was used in our previous study [15].
-This dataset combines the data from the greedy agent and the N-1 agent.
-The data from the N-1 agent on the full network is used.
-The data from the greedy agent on a subset of the N-1 networks is used.
-These are the N-1 networks with lines 0, 2, 4, 5, 6, and 12 disabled, which the greedy agent can operate well [15].
+We use the dataset generated by the greedy and N-1 expert agents, described in our previous work deJong2024 .
+Both agents use extensive load flow simulations to find the optimal action.
+The greedy agent selects actions by minimizing the maximum line loading of the grid in its current state, whereas the N-1 agent minimizes the maximum line loading under the disablement of any power line.
+The N-1 agent demonstrated superior performance compared to the greedy agent in settings both with and without outages.
+Both agents use an activity threshold below which “do-nothing” actions are automatically selected.
+State-action pairs from such do-nothing actions, or from episodes that the expert agents could not successfully complete, were excluded from the dataset.
+Do-nothing actions selected above the activity threshold were included.
+We constructed two datasets.
+First, an in-distribution (ID) dataset, containing network configurations on which the models are trained and evaluated.
+This dataset was used in our previous study deJong2024 .
+This dataset incorporates state-action pairs from both the greedy agent and the N-1 agent.
+Specifically, it incorporates state-action pairs from the N-1 agent operating on the full network, and from the greedy agent operating on a subset of N-1 networks (lines 0, 2, 4, 5, 6, and 12 disabled).
+These specific N-1 networks were selected because the greedy agent can operate them well deJong2024 .
 Second, an out-of-distribution (OOD) dataset, which is used to investigate generalization. This dataset comprises data from the greedy agent on the N-1 networks with lines 1 and 3 disabled.
-On both datasets, we split the datapoints into 70/10/20 train/validation/test sets based on their scenario.
-The sizes of the partitions are listed in Table 3.
-In this paper, ID or OOD networks refer to the networks in either the ID or OOD datasets.
-ID or OOD outages refer to the outages present in either ID or OOD networks.
+For both datasets, we split the datapoints into 70/10/20 train/validation/test sets based on scenario.
+The partition sizes are listed in Table 2.
+In this paper, when we refer to “ID networks” or “OOD networks”, we refer to the network variations in the ID or OOD datasets, respectively.
+Likewise, when we refer to “ID outages” or “OOD outages”, we refer to the outages present in respectively ID networks or OOD networks.
 [t]    Train    Val    Test    ID    196,477    26,228    59,150    OOD    41,788a    5,690a    12,724 a These datapoints are used to train a model to estimate an upper bound to generalization (see Sec.
-4.8).
-Table 3: The number of datapoints in the datasets and train/val/test sets.
+3.8).
+Table 2: The number of datapoints in the datasets and train/validation/test sets.
 3.2 Datapoints
-Each datapoint includes the features per object, the topology vector, and the expert action.
-The features for each object type are listed in Table 4.
-The object features are normalized.
+Each datapoint includes each object’s features, the topology vector, and the expert action.
+The features per object type are listed in Table 3.
+Object features are normalized.
 If a line is disabled, its endpoint features are zero-imputed, and the corresponding topology vector values are -1.
@@
 3.3 FCNN
-The fully connected neural network (FCNN) consists of an input layer, multiple hidden layers, and an output layer.
-The input vector is obtained by flattening the input features into a vector and concatenating the topology vector.
+The fully connected neural network (FCNN) model consists of an input layer, multiple hidden layers, and an output layer.
+The input vector is constructed by flattening the object features into a single vector and concatenating the topology vector.
 The hidden layers use the ReLU activation function.
@@
 3.4 Homogeneous GNN
-Applying a graph neural network to the power grid requires a graph representation of the power grid.
-We represent the grid objects as nodes.
-We connect nodes if the corresponding objects are attached to the same busbar or are corresponding line endpoints.
+Applying a graph neural network to the power grid requires a graph representation of the grid.
+We represent the grid objects as nodes within the graph.
+Edges are set between nodes if the corresponding objects are connected to the same busbar or represent (corresponding) line endpoints.
 We call this graph representation homogeneous as it does not consider edge types.
 We call the associated GNN the HomGNN.
-Figures 2a and 2b display an example grid and its homogeneous representation.
+Figures 2a and 2b respectively display an example grid and its homogeneous graph representation.
 (a) Example Grid
@@
 Different types of grid objects have different features (see Table 3). We use multiple two-layer perceptrons to embed the varying object features into a common embedding:
-$\textbf{h}_{u,0}=\begin{cases}\text{MLP}_{gen}(\textbf{x}_{u})&\text{node $u$ represents a generator}\\ \text{MLP}_{load}(\textbf{x}_{u})&\text{node $u$ represents a load}\\ \text{MLP}_{line}(\textbf{x}_{u})&\text{node $u$ represents a line endpoint},\end{cases}$
+[…]
 where $\textbf{h}_{u,0}$ and $\textbf{x}_{u}$ are, respectively, the initial node embedding and the features of node $u$.
-The embeddings of subsequent layers are computed with the message passing rule [36]:
+The embeddings of subsequent layers are computed with the message passing rule 4700287 :
 $\textbf{h}_{u,k+1}=\sigma(W_{self,k}\textbf{h}_{u,k}+W_{neighbor,k}\sum_{v\in\mathcal{N}(u)}\textbf{h}_{v,k}+\textbf{b}_{k}),$
@@
 3.5 Busbar Information Asymmetry
-The aforementioned graph neural network formulation has a problem.
-Each node’s output specifies whether to switch the corresponding object between busbars.
+The aforementioned graph neural network formulation suffers from a problem.
+The output of each node specifies whether to switch the corresponding object across busbars.
 However, messages from objects on the current busbar are passed, while messages from the objects on the other busbar are not.
 We call this the busbar information asymmetry.
-This is intuitively problematic as a switching decision requires information about the other busbar.
-This also decreases the model’s expressiveness.
-Objects at the same substation but on different busbars and objects at entirely different substations cannot be distinguished.
-Moreover, the absence of inter-busbar connections also decreases the graph’s connectivity.
+We posit that information about objects on both busbars should be equally available, as both groups of objects are involved in actions.
+Moreover, this decreases the model’s expressiveness: objects at the same substation but on different busbars and objects at entirely different substations cannot be distinguished.
+Lastly, the absence of inter-busbar connections also decreases the graph’s connectivity.
 This can result in longer paths that may require deeper GNNs.
 3.6 Heterogeneous GNN
-Our solution is to represent the grid with a heterogeneous graph representation.
-The associated GNN is called the HetGNN.
-This heterogeneous representation features three edge types: one for objects on the same busbar, one for objects on the other busbar (but at the same substation), and one for corresponding line endpoints.
-This is displayed in Figure 2c.
+To address this problem, we propose a heterogeneous graph representation, and a corresponding GNN, the HetGNN This heterogeneous representation features three edge types: edges connecting objects on the same busbar, edges connecting objects on the other busbar but within same substation, and edges connecting corresponding line endpoints.
+The heterogeneous graph representation is displayed in Figure 2c.
 The message passing rule is adapted to consider a weight and neighborhood aggregation per edge type:
@@
 The $W_{line,k}\textbf{h}_{line,k}$ term represents the addition of the embedding of a node connected by a power line.
-It is therefore omitted for nodes that do not represent line endpoints.
+This term is therefore omitted for nodes that do not represent line endpoints.
 Figure 2e displays an example of heterogeneous message passing.
@@
 $L=$ $\textrm{mean}(-\textbf{w}(\textbf{y}\log(\textbf{p})+(\textbf{1}-\textbf{y})\log(\textbf{1}-\textbf{p}))),$
+where $\textbf{w},\textbf{y}$, and p denote the label weight, target, and prediction vectors, respectively.
+We introduce label weights to prevent the prediction vectors from collapsing to zeros.
+This occurs as the majority of target values are zero.
+A lower label weight $0<\alpha<1$ is assigned to labels that do not correspond to objects at either the target or predicted substation:
 […]
-$_{i}=\begin{cases}1&\text{object $i$ corresponds to the target substation}\\ 1&\text{object $i$ corresponds to the predicted substation}\\ \alpha&\text{otherwise}\end{cases}\quad\forall w_{i}\in\textbf{w}.$ w .
 There is no target substation if the target action is a do-nothing action.
-The predicted action is classified as a do-nothing action, without a predicted substation, if all predictions $_{i}\in\textbf{p}$ p do not exceed $0.5$.
+The predicted action is classified as a do-nothing action, without a predicted substation, if all predictions $p_{i}\in\textbf{p}$ do not exceed $0.5$.
 Otherwise, the predicted substation is the substation where the predictions $\textbf{p}_{s}$ at that substation $s\in S$ maximize
@@
 However, not every prediction vector corresponds to an action in the filtered action space (see Sec.
-3.3).
+2.3).
 We apply a postprocessing step that replaces the model’s prediction p with the nearest action.
@@
 d The ReLU negative slope parameter was not tuned.
-Figure 3: The training curves of the five models per model type. Green lines show the HetGNNs, yellow lines the FCNNs, and blue lines the HomGNNs.
+Figure 3: The training curves of the five models per model type.
 3.9 Imitation Learning Agents
@@
 The verify+N-1 agent functions similarly but with the N-1 agent.
-Each agent also uses the activity threshold used by the expert agents [15].
+Each agent also uses the activity threshold used by the expert agents deJong2024 .
 4 Results
@@
 a This lists the accuracy without the postprocessing step described in Sec.
-4.7.
+3.7.
 b These datapoints represent the default topology, i.e., without split substations.
 c These datapoints represent non-default topologies, i.e., with split substations.
-Table 6 lists the accuracies of the different models on the various data partitions and the ID/OOD network groups.
-Figure 4 shows the test accuracies of the models on both datasets.
-On the ID dataset (rows 1 to 3), the accuracies remain limited to approximately 80%.
+Table 5 presents the accuracies of the different models across the data partitions and the ID/OOD network groups.
+Figure 4 displays the test accuracies of the models on both datasets.
+On the ID dataset (rows 1-3), the accuracies remain limited to approximately 80%.
 On each split, the HetGNN achieves the highest accuracy, followed by the FCNN and, lastly, the HomGNN.
-Each model type shows a tiny drop in accuracy between the train, validation, and test sets, indicating slight overfitting.
-Row 4 displays that accuracy drops slightly without the postprocessing step described in Sec.
-4.7.
-The default topology appears frequently in the ID dataset, comprising 46% of the test set.
-In this topology, all objects are attached to the same busbar.
-This avoids the busbar asymmetry problem and consequently negates the theoretical advantage of the HetGNN over the HomGNN.
+A minor decrease in accuracy is observed between the training, validation, and testing sets for all model types, suggesting a slight degree of overfitting.
+As shown in row 4, accuracy decreases marginally when the post-processing step outlined in Section 3.7 is omitted.
+The default topology, characterized by all objects connected to the same busbar, appears frequently within the ID dataset, constituting 46% of the test set.
+This configuration avoids the busbar asymmetry problem, thereby negating the theoretical advantage of the HetGNN over the HomGNN.
 As shown in row 5, the accuracies on the standard topology are relatively similar.
-As shown in row 6, the accuracies vary considerably between topologies with split busbars.
+Conversely, as shown in row 6, the accuracies vary considerably in topologies with split busbars.
 Figure 4: The test accuracies of the model types on the ID dataset set (left) and OOD dataset (right). The ranges indicate the maximum and minimum accuracy.
-The GNNs obtain far higher accuracies on the OOD dataset than the FCNNs (rows 7 to 9).
-The HetGNNs also obtain higher accuracies than the HomGNNs.
-However, the performance of either the HomGNN or the HetGNN on the OOD dataset is still substantially lower than models trained on OOD data (column ’OOD-GNN’).
+The GNNs achieve far higher accuracies on the OOD dataset than the FCNNs (rows 7-9).
+The HetGNNs also achieve higher accuracies than the HomGNNs.
+However, the performance of either the HomGNN or the HetGNN on the OOD dataset is still substantially lower than models specifically trained on OOD data (column ’OOD-GNN’).
 4.2 Error Analysis
-We repeated the error analysis previously performed on the FCNN [15] to the two types of GNN models.
-We again find that the class imbalance and class overlap play a role in the limited accuracy.
-Figure 5 shows that both GNN types predict infrequent classes disproportionately infrequently.
-We investigated pairs of classes that were often confused, i.e., pairs of classes that were often mistaken for one another in prediction.
-Figure 6 shows that the errors of frequently confused classes are in overlapping regions.
+We repeated the error analysis previously performed on the FCNN deJong2024 for both GNN model types.
+As with the FCNN, class imbalance and overlap contribute to the limited accuracy.
+As illustrated in Figure 5, both GNN types exhibit a tendency to under-predict infrequent classes.
+We investigated frequently confused class pairs, i.e.
+classes commonly mistaken for each other during prediction.
+Figure 6 indicates that the errors associated with these frequently confused classes occur within overlapping regions of the feature space.
 Figure 5: The log-distributions of the classes in the ID validation set.
@@
 This finding is consistent among the models and confused classes.
-Inspection of the nearest neighbors supports the notion that the class overlap contributes to low accuracy.
-We applied one of the HetGNNs to a subset of 2,500 validation data points from the full network.
-The accuracy over the data points whose nearest neighbor is in the same class is 92.97%, but only 44.44% for those whose nearest neighbor is in another class.
-Other models showed similar results.
+Analysis of nearest neighbors further supports the hypothesis that class overlap contributes to reduced accuracy.
+We applied one of the HetGNNs to a subset of 2,500 validation datapoints.
+For data points whose nearest neighbor belongs to the same class, the model achieves an accuracy of 92.97%.
+In contrast, for data points whose nearest neighbor belongs to a different class, the accuracy drops to 44.44%.
+Similar trends were observed across models.
 4.3 Graph Smoothness
-The tendency of graph neural networks to suffer from oversmoothing is well-documented [3, 23].
-We compute the mean average distance (MAD) values to measure graph smoothness [3].
-MAD values quantify the similarity of node embeddings.
-Small MAD values indicate that node embeddings are similar and, thus, that the graph is smooth.
-The MAD values are computed between neighbor nodes over the first 1000 validation datapoints.
-Figure 5.4 shows the MAD values between the various layers.
-As visible, the MAD values in the middle GNN layers are consistently higher in the HetGNN than in the HomGNN.
-This suggests that the heterogeneous representation leads to a less smooth local embedding.
-This finding is consistent among the models.
+Given the well-documented tendency of graph neural networks to exhibit oversmoothing chen_oversmoothing , 10.5555/3504035.3504468 , which can degrade model performance, we assess smoothness of the graphs produced by both GNNs.
+We calculate the mean average distance (MAD) values, which quantify the distance between neighboring node embeddings.
+Lower MAD values correspond to higher embedding similarity and thus increased graph smoothness.
+Figure 4.4 shows the MAD values over the various layers for both GNN types, calculated on the first thousand validation datapoints.
+As shown, the MAD values of the middle GNN layers are consistently higher in the HetGNN than in the HomGNN.
+This finding is consistent across the different models per type, suggesting that oversmoothing is a smaller concern for the HetGNN.
 4.4 Graph Diameter
@@
 We measure this by the diameter, i.e., the shortest path length between the two most distant nodes.
-We calculate the diameter by finding the lowest exponent of the adjacency matrix that produces a matrix without zero entries [7].
+We calculate the diameter by finding the lowest exponent of the adjacency matrix that produces a matrix without zero entries Duncan2004PowersOT .
 The diameters of the 25 most common topologies are shown in Figure 4.4.
@@
 Table 6: The percentage of days completed by agents in various regimes. The agents were evaluated on the test scenarios.
-Table 7 shows the days completed by the various agents in the different settings.
-The ’Full Network’ column describes the setting without outages.
-The ’ID Outages’ describes a setting with an opponent that causes random outages present in the ID dataset (see Sec.
-4.1).
-The ’OOD Outages’ describes a setting with an opponent that causes random outages present in the OOD dataset.
-The agents can be ordered by performance: Do-Nothing $<$ Naive $<$ Verify $<$ Greedy Hybrid $<$ N-1 Hybrid.
-The naive agents (rows 4 to 6) and verify agents (rows 8 to 10) come close to the performance of the N-1 expert agents (row 3).
-The hybrid agents (rows 12 to 14, and 16 to 18) can match the N-1 expert agents.
-On the setting with no or ID outages (columns ’Full Network’ and ’ID Outages’) the HetGNN network consistently performs best, followed by the FCNN and, finally, the HomGNN (rows 4 to 6, and 8 to 10).
-Additional simulation reduces or nullifies the differences (rows 12 to 14, and 16 to 18).
-On the setting with OOD outages (column ’OOD Outages’), the GNNs outperform the FCNN (rows 4 to 6, and 8 to 10).
-However, additional simulation also decreases this effect (rows 12 to 14, and 16 to 18).
-Furthermore, the performance of the models trained on the ID dataset remains substantially lower than models trained on OOD dataset (rows 7 and 11), even with additional simulation (rows 15 and 19).
-5.6 Efficiency
-Finally, we consider the inference speed of the different agents.
-The inference speed was measured on an Apple M1 CPU with minimal background processes on the first fifty validation scenarios.
-Table 5.6 lists the durations per model type and agent.
-Figure 5.6 plots these values against the agent’s performance on the full-network topology.
+Table 6 shows the percentage of days completed per agents across the different settings.
+The ‘Full Network’ column describes the setting without outages.
+The ‘ID Outages’ and ‘OOD Outages’ columns correspond to settings with random outages derived from the in-distribution and out-of-distribution datasets, respectively (see Sec.
+3.1).
+Agent performance follows a consistent order: Do-Nothing << Naive << Verify << Greedy Hybrid << N-1 Hybrid.
+The naive agents (rows 4-6) and verify agents (rows 8-10) achieve performance near to the N-1 expert agent (row 3).
+The hybrid agents (rows 12-14 and 16-18) can match the performance of the N-1 expert agent.
+On the setting with no or ID outages (columns ’Full Network’ and ’ID Outages’) the HetGNN network consistently performs best, followed by the FCNN and, lastly, the HomGNN (rows 4-6 and 8-10).
+Additional simulation reduces or nullifies the differences (rows 12-14 and 16-18).
+On the setting with OOD outages (column ’OOD Outages’), the GNNs outperform the FCNN (rows 4-6 and 8-10).
+However, additional simulation also reduces this effect (rows 12-14 and 16-18).
+Furthermore, the performance of the models trained on the ID dataset remains substantially lower than models directly trained on OOD dataset (rows 7 and 11), even with additional simulation (rows 15 and 19).
+4.6 Computational Efficiency
+We finally consider the inference speed of the different agents.
+Inference times are measured using an Apple M1 CPU with minimal background processes across the first fifty validation scenarios.
+Table 4.6 lists the mean duration per model and agent type.
+Figure 4.6 plots these values against the agent’s performance on the setting with OOD outages.
 The GNNs, particularly the HetGNNs, are considerably slower than the FCNNs.
-However, these differences are small relative to the simulation times of the expert agents, which are orders of magnitude higher.
+However, their computation times remain orders of magnitudes smaller than those of the expert agents.
 Most importantly, all GNN agents have a favorable combination of speed and performance.
@@
 […]
-tableInference durations of various agents.
-The hybrid models all use the FCNN.
-The differences between model types for the hybrid agents are relatively insignificant.
-Agent Duration ($\upmu$s) Greedy 6.77E5 N-1 4.70E6 Naive FCNN 5.74E2 Naive HomGNN 2.81E3 Naive HetGNN 3.84E3 Verify FCNN 1.14E4 Verify HomGNN 1.50E4 Verify HetGNN 1.58E4
+[…]
 5 Discussion
-The heterogeneous GNN consistently performs best, followed by the FCNN, and lastly, the homogeneous GNN.
-This is true for both the accuracy and grid operation.
-The superiority of the HetGNN to the FCNN indicates that GNNs can exploit the graph structure better.
-The inferiority of the HomGNN can be attributed to the busbar asymmetry problem.
-However, smoother graphs and longer path lengths might also play a role.
-The HetGNN’s coarser graphs can be attributed to the more diverse edge representation.
-The GNNs take considerably longer to evaluate than the FCNNs, although this difference is very tiny compared to the expert agents.
-The GNNs achieve far higher accuracy than the FCNNs on the OOD networks.
-This is also reflected in the simulation performance, although the effect is smaller.
-Although GNNs show a superior ability to generalize to OOD networks, they still perform considerably worse than GNNs trained on these networks.
-There seemed to be a limit to the accuracy of the trained models.
-This can be attributed to class imbalance and class overlap.
-We hypothesize that the class overlap originates from the Grid2Op simulate function.
-It is possible that identical network states have diverging forecasts and consequently diverging actions.
-The limited accuracy does not stop the ML models from successfully completing many of the grid operation days.
-The naive ML models come close to the performance of the expert agents.
-The hybrid agents can match them.
-Hybrid agents can outperform expert agents in specific settings, which is surprising.
-We hypothesize that this results from omitting the state-action pairs from failed runs.
-The model bias towards frequent actions might also result in more robust topologies.
-An interesting effect is that model differences are more pronounced in accuracy than simulation performance.
-Similarly, the limited accuracy does not stop the models from operating the grid well.
-We have two explanations for this.
-Firstly, a misprediction does not necessarily lead to a game-over.
-This is probably particularly true in regions with class overlap.
-Secondly, difficult days led to the generation of many relatively uncommon datapoints.
-Failure to predict many such datapoints would consequently affect the completion of only few days.
-7 Future Work
-Our findings indicate the promise of fully graphical neural networks for topology control methods.
-However, further research is necessary to establish this.
-Future work should compare fully GNN models with models that only use GNNs for feature extraction.
-Similarly, the heterogeneous graph representation should be introduced into other approaches.
-Moreso, it would be good to apply these methods in a reinforcement learning framework.
-Further efforts are likely also necessary to scale this approach to larger grids.
-Fully GNN models might require very deep GNNs to model large graphs.
-Very deep GNNs are associated with additional challenges [46].
-Future research could focus on methods to downsample grid graph representations, e.g., through trainable graph pooling [12].
-Alternatively, it might be beneficial to combine GNN layers and other layers to capture long-range patterns.
-The imitation learning approach used here can also be developed further.
-Our approach suffers from distribution shift, a compounding divergence between the behavior of the ML and the expert agent [1].
-More advanced IL frameworks, such as DAgger, address this [34].
-Investigating other expert agents and developing an improved dataset would also be valuable.
-Finally, the generalization capabilities of graph neural networks should be studied more extensively.
-Future studies should investigate how factors such as the diversity of training data and differences between the ID and OOD data affect generalization.
+5.1 Interpretation
+The heterogeneous GNN consistently achieves the highest performance, in regards to both classification accuracy and grid operation ability.
+The homogeneous GNN exhibits the lowest performance, being surpassed by the FCNN.
+The superior performance of the HetGNN compared to the FCNN demonstrates the ability of GNNs to effectively exploit graph structure.
+The inferior performance of the HomGNN can be attributed to the busbar asymmetry problem.
+Smoother graphs and longer path lengths might further contribute the inferior performance.
+The more diverse edge representation of the HetGNN likely reduces the oversmoothing effect of GNNs.
+Concerning efficiency, the GNNs take considerably longer to evaluate than the FCNNs.
+However, this difference is negligible in comparison with the expert agents.
+A question that naturally arises is why homogeneous graph neural networks have exhibited good performance in related research yoon2020winning .
+We believe that the answer lies in the fact that models used in other research combine GNN layers with other layers (e.g., linear layers), which mitigates the busbar asymmetry problem.
+This presents an interesting topic for future work.
+GNNs demonstrate a superior ability to generalize to out-of-distribution (OOD) network configurations than FCNNs.
+This effect is present in both accuracy and, albeit with reduced magnitude, simulation performance.
+Despite their superior generalizability, GNN OOD performance remains substantially below that of models directly trained on OOD network configurations.
+The accuracy of the trained models appears to reach a plateau, which may be attributed to issues such as class imbalance and class overlap.
+We hypothesize that the observed class overlap originates from the Grid2Op simulate function.
+It is possible that identical network states have diverging forecasts and consequently diverging actions, causing class overlap.
+This limitation in classification accuracy does not preclude the models from achieving good operation performance, however.
+The naive ML agents achieve operational performance close to the expert agents, whereas the hybrid agents can match them.
+Counterintuitively, hybrid agents sometimes achieve performance superior to the expert agents.
+A potential source for this effect is the exclusion of samples from the dataset for days that the expert agents could not successfully complete, positively biasing the models to successful behavior.
+The model bias towards frequent actions might also result in robust topologies that can accommodate a wider set of injection profiles.
+An interesting effect is that model differences are more pronounced with respect to accuracy than operational performance.
+Similarly, the limited accuracy does not stop the models from effectively operating the grid.
+We propose two contributing factors.
+First, a misprediction does not necessarily lead to a game-over; the class overlap suggests that there are often multiple viable actions.
+Second, certain challenging days generate a high volume of uncommon data points.
+Misclassification of these datapoints affects a considerable subset of the dataset but is limited to only few operational days.
+5.2 Future Work
+We propose the integration of the methods developed here, particularly the heterogeneous graph representation, into other approaches as a focus of future research.
+The heterogeneous graph representation could be integrated seamlessly into previously proposed model architectures.
+It is particularly still an open question how HetGNN layers interact with non-graphical model layers, warranting investigation.
+Incorporating the method into a reinforcement learning framework is also a crucial next step.
+It is particularly promising to integrate these approaches into a multi-agent reinforcement learning (MARL) framework, which has recently gained attention for topology control de_Mol_2025 , vandersar2023 .
+Within MARL, the combinatorially large action space can be factorized across sub-agents.
+The agents developed here could be deployed as localized sub-agents within such a system, offering a natural path toward scaling to larger grids.
+Although existing approaches have applied imitation learning only in it’s simplest form, there are compelling reasons to apply more advanced methods.
+Notably, the simple behavioral cloning learning setup suffers from distribution shift, a compounding divergence between the behavior of the ML and expert agent belkhale2024data .
+DAgger is a more advanced method that addresses this problem pmlr-v15-ross11a .
+Additionally, it would also be valuable to further explore expert agents and develop better datasets.
+Finally, given the dynamic nature of power grids, generalization of machine learning models remains a topic to be studied more extensively.
+There is yet little knowledge about how factors such as the diversity of training data and differences between ID and OOD data affect generalization.
+Data augmentation might be a promising technique to obtain models with wide applicability.
+6 Conclusion
+To guarantee energy security, transmission system operators (TSOs) must effectively manage grid congestion.
+Topology control guided by machine learning presents a powerful tool for TSOs to do so.
+While graph neural networks (GNNs) are well-suited to model the grid’s structure, we identify that their straightforward application suffers from a critical limitation: the busbar asymmetry problem.
+To address this, we introduce a heterogeneous graph representation that resolves the asymmetry and enables effective learning.
+Heterogeneous graph neural networks outperform homogeneous graph neural network, the latter of which fails to outperform the fully-connected neural network (FCNN) baseline.
+In turn, heterogeneous GNNs also surpass fully connected neural networks (FCNNs), demonstrating the value of incorporating grid topology explicitly.
+This performance gap is most pronounced in out-of-distribution settings, where FCNN performance deteriorates significantly, while GNNs retain a higher degree of generalization.
+Importantly, agents trained with these methods achieve operational performance comparable to expert systems, while offering greatly improved computational efficiency.
+These findings demonstrate the potential of GNN-based agents as a solution to support TSOs in real-time grid management.
 7 Statements
@@
 7.5 Code & Data Availability
-Our code \@footnotemark and data [16] are publicly available.
+Our code 2 and data deJong2024imitation are publicly available.
+7.6 Declaration of generative AI and AI-assisted technologies in the writing process.
+During the preparation of this work the author(s) used Google Gemini in order to increase the clarity and readability of the text.
+After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
+The tool was only used for the reviewing and rephrasing of existing text, not the generation of new text.